# FS-Net: Full Scale Network and Adaptive Threshold for Improving Extraction of Micro-Retinal Vessel Structures

### v3

@@ 5.2 Performance Evaluation
The performance of this work is compared with the state-of-the-art studies including U-Net ronneberger2015u , Attention U-Net oktay2018attention , U-Net++ zhou2018unet++ , U-Net3+ huang2020unet , R2U-Net alom2019recurrent , SGL zhou2021study , RV-GAN kamran2021rv , FR-UNet liu2022full . We have trained and evaluated the first four U-Net versions, and for the rest of the works, the performance reported by the authors is taken for comparison.
@@
We trained and evaluated U-Net variants under the same conditions as our model, incorporating Gamma correction and CLAHE jiang2019automatic on grayscale images, significantly improving fundus image clarity. Our U-Net variants outperformed those in existing research.
@@ 5.3 Ablation Study
Table 3: Ablation study of the proposed work on DRIVE dataset
Methods
Sen.
Spe.
F1
Acc.
AUC
IoU
BL
0.7827
0.9887
0.8218
0.9704
0.9877
0.6981
BL+EB
0.7894
0.9882
0.8236
0.9706
0.9878
0.7007
BL+EB+BE
0.7929
0.9878
0.824
0.9705
0.988
0.7012
BL+EB+BE+SE
0.7901
0.9886
0.8259
0.9709
0.9884
0.7038
BL+EB+BE+SE+AT
0.8421
0.9828
0.8314
0.9702
0.9884 0.7116
@@ Methods
Table 4: Performance of cross-training evaluation
Train
Test
Methods
Sen.
Spe.
Acc.
AUC
DRIVE STARE
HAnet wang2020hard
0.8187
0.9699
0.9543
0.9648
RE-GAN zhou2021refined
0.8334
0.9764
0.9613
0.9718
MFI-Net ye2022mfi
0.7805
0.9741
0.9550
0.9747
FS-Net 0.892
0.9712
0.9651 0.9871
STARE DRIVE
HAnet wang2020hard
0.7140
0.9879
0.9530
0.9758
RE-GAN zhou2021refined
0.7412
0.9830
0.9519
0.9643
MFI-Net ye2022mfi
0.7313
0.9867
0.9538
0.9762
FS-Net
0.7402
0.9889 0.967 0.9793
@@ 5.4 Cross Training Evaluation
The ultimate goal of any deep learning solution is to create a system that is capable of high generalization. The situation is different in the case of retinal vascular segmentation. There are three frequently used datasets, namely DRIVE, CHASE-DB1, and STARE, and all of the works presented so far have been dedicated to training and testing in a single dataset, despite the fact that all of these datasets are retinal datasets. However, to take this study to the next level, it must be able to generalize well regardless of dataset type, as in the case of real-world scenarios. There have been a number of studies on cross-training evaluation: training in one dataset and testing in another wang2020hard ; zhou2021refined ; ye2022mfi . In this work, we ran cross-training analyses on the DRIVE and STARE datasets and obtained impressive results when compared to previous studies. When trained on the DRIVE dataset and tested on the STARE dataset, our model exhibited very high sensitivity, accuracy, and AUC, owing to its superior accuracy in segmenting multi-scale vascular structures. When the model is trained on STARE and evaluated on DRIVE, it achieves better results in all metrics except sensitivity. This is because the STARE dataset does not have as many small-scale and branch vessels as DRIVE, but it is still just 0.001 behind RE-GAN.
@@ 6 Conclusion
In this research, we contributed to the main challenging problem of microvascular extraction in retinal vessel segmentation. The proposed solutions have different components, such as encoder boosters to keep track of information extracted by the encoder architecture, bottleneck enhancement blocks where full-scale features are learned better, squeeze and excitation blocks where feature representation of the network is improved by capturing channel dependencies, sigmoid smoothing, and adaptive threshold algorithms where microvascular segments are extracted at a high rate. The significance of all these components was evaluated by conducting an ablation study, and the results obtained demonstrated their effectiveness for the application at hand. Further, the proposed solution was also passed through cross-training evaluation in order to ascertain how well our model, when trained on one dataset, can generalize when it is tested on another dataset, and as a result, good performance was recorded. We also intend to investigate the efficacy of the proposed approach using multi-modal data in future work.
7 Acknowledgements.
Part of this work was supported by the Skolkovo Institute of Science and Technology - University of Sharjah Joint Projects: Artificial Intelligence for Life.
Declaration of Competing Interest

### v4

@@ 5.2 Performance Evaluation
The performance of this work is compared with the state-of-the-art studies including U-Net [5], Attention U-Net [36], U-Net++ [37], U-Net3+ [29], R2U-Net [39], SGL [22], RV-GAN [24], FR-UNet [23]. We have trained and evaluated the first four U-Net versions, and for the rest of the works, the performance reported by the authors is taken for comparison.
@@
We trained and evaluated U-Net variants under the same conditions as our model, incorporating Gamma correction and CLAHE [40] on grayscale images, significantly improving fundus image clarity. Our U-Net variants outperformed those in existing research.
@@
To demonstrate the adaptability of our model across different conditions, we tested our work on another medical segmentation dataset known as DCA1. This dataset comprises X-Ray coronary angiograms characterized by comparatively low-resolution grayscale images. Since the dataset does not come with an official training and testing split, researchers must randomly partition it for their analyses. Consequently, this lack of a standardized split makes it challenging to compare the performance of our model with other studies. To facilitate a valid comparison, we trained four different models from scratch using the same dataset configuration. Our model achieved superior performance compared to other works, as shown in Table 2.
The proposed approach also has advantages regarding computational complexity compared to previous works. Although the model is not considered to be very lightweight, which is not our primary concern, the size of the parameters is acceptable considering the complexity of integrated architectures. The number of parameters and the floating-point operations for the U-Net architecture, its variants, and our model are presented in Table 3.
Table 3: Complexity comparison between different models, where the first FLOPs column corresponds to an input size of 1$\times$512$\times$512 and the second column corresponds to an input size of 1$\times$48$\times$48.
Methods Params (M) FLOPs (G) FLOPs (G)
U-Net[5] 31.04 218.52 1.92
Attention U-Net [36] 34.88 266.15 2.34
U-Net++[37] 36.63 551.95 4.85
U-Net3+[29] 26.97 791.56 6.96
FS-Net 7.14 216.95 1.9
@@ 5.3 Ablation Study
Table 4: Ablation study of the proposed work on DRIVE dataset.
Methods Sen. Spe. F1 Acc. AUC IoU
BL 0.7827 0.9887 0.8218 0.9704 0.9877 0.6981
BL+EB 0.7894 0.9882 0.8236 0.9706 0.9878 0.7007
BL+EB+BE 0.7929 0.9878 0.824 0.9705 0.988 0.7012
BL+EB+BE+SE 0.7901 0.9886 0.8259 0.9709 0.9884 0.7038
BL+EB+BE+SE+AT 0.8421 0.9828 0.8314 0.9702 0.9884 0.7116
@@
Table 5: Performance of cross-training evaluation.
Train Test Methods Sen. Spe. Acc. AUC
DRIVE STARE HAnet [19] 0.8187 0.9699 0.9543 0.9648
RE-GAN [41] 0.8334 0.9764 0.9613 0.9718
MFI-Net [42] 0.7805 0.9741 0.9550 0.9747
FS-Net 0.892 0.9712 0.9651 0.9871
STARE DRIVE HAnet [19] 0.7140 0.9879 0.9530 0.9758
RE-GAN [41] 0.7412 0.9830 0.9519 0.9643
MFI-Net [42] 0.7313 0.9867 0.9538 0.9762
FS-Net 0.7402 0.9889 0.967 0.9793
@@ 5.4 Cross Training Evaluation
The ultimate goal of any deep learning solution is to create a system that is capable of high generalization. The situation is different in the case of retinal vascular segmentation. There are three frequently used datasets, namely DRIVE, CHASE-DB1, and STARE, and all of the works presented so far have been dedicated to training and testing in a single dataset, despite the fact that all of these datasets are retinal datasets. However, to take this study to the next level, it must be able to generalize well regardless of dataset type, as in the case of real-world scenarios. There have been a number of studies on cross-training evaluation: training in one dataset and testing in another [19, 41, 42]. In this work, we ran cross-training analyses on the DRIVE and STARE datasets and obtained impressive results when compared to previous studies. When trained on the DRIVE dataset and tested on the STARE dataset, our model exhibited very high sensitivity, accuracy, and AUC, owing to its superior accuracy in segmenting multi-scale vascular structures. When the model is trained on STARE and evaluated on DRIVE, it achieves better results in all metrics except sensitivity. This is because the STARE dataset does not have as many small-scale and branch vessels as DRIVE, but it is still just 0.001 behind RE-GAN.
@@ 6 Conclusion
In this research, we contributed to the main challenging problem of microvascular extraction in retinal vessel segmentation. The proposed solutions have different components, such as encoder boosters to keep track of information extracted by the encoder architecture, bottleneck enhancement blocks where full-scale features are learned better, squeeze and excitation blocks where feature representation of the network is improved by capturing channel dependencies, sigmoid smoothing, and adaptive threshold algorithms where microvascular segments are extracted at a high rate. The significance of all these components was evaluated by conducting an ablation study, and the results obtained demonstrated their effectiveness for the application at hand. Further, the proposed solution was also passed through cross-training evaluation in order to ascertain how well our model, when trained on one dataset, can generalize when it is tested on another dataset, and as a result, good performance was recorded. Future work will focus on addressing the three-fold generalization challenges that emerged during the course of our research: visual and degradation style shifts, diagnostic pattern diversity, and data imbalance. These issues are prevalent in complex real-world scenarios, such as microvascular extraction in retinal vessel segmentation. To tackle these challenges, we aim to develop a novel unified and generalizable framework that can effectively adapt to diverse clinical environments and data variations. Further enhancements of our model will integrate advanced techniques for domain adaptation, robust feature learning, and new data augmentation strategies to mitigate the impact of these generalization issues.
7 Acknowledgements
Part of this work was supported by the Skolkovo Institute of Science and Technology - University of Sharjah Joint Projects: Artificial Intelligence for Life, project ”Towards an Explainable Diabetic Retinopathy Grading Model”.
Declaration of Competing Interest.